OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 I haven't done these type of problems in a while, so I forgot how to do them. If someone could help me, that would be much appreciated. Thank you!

OpenStudy (anonymous):

do you know how to make the equation equal to y

OpenStudy (anonymous):

@mcclm666 you still here

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

5x-3y=-6 -5x -5x -3y=-6-5x divide both sides by -3 y=(5/3)x+2 so the slope is the number that is multiplied by x in this case(5/3) the y-intercept is the last number of the equation in this case 2

OpenStudy (anonymous):

Okay thank you so much! x